# Michel F. C. Haddad

### v1

@@ 6 Simulation study
In our simulations, we set the number of covariates to $p=100$ and the $j$th element in the coefficient vector $\beta$ to $0.4/i^{2}$ for $j\in\{1,...,p\}$, which implies a quadratic decay of covariate importance in terms of confounding. We consider two sample sizes of $n=2000,4000$. DiD estimation is based on the sample analog of equation (2). In our simulation design, there are only two periods $t$ such that $T=t$ and $T=t-1$ in equation (2) correspond to $T=1$ and $T=0$, respectively. Furthermore, there is no history of previous treatments in $T=0$ such that $\mathbf{D_{t-1}}=\emptyset$ in our simulations. However, from an econometric perspective, there is no difference between controlling for the treatment history $\mathbf{D_{t-1}}$ or covariates $X$, which are both time-varying sets of confounders. For this reason, not considering $\mathbf{D_{t-1}}$ in our simulations comes without loss of generality as we could easily relabel some elements in $X$ to represent $\mathbf{D_{t-1}}$.
@@
For estimation, we apply the didcontDML function of the causalweight package by Bodory and Huber (2018) for the statistical software R (R Core Team (2020)). We run this DML-based DiD estimator for repeated cross-sections with 3-fold cross-fitting and estimate the nuisance parameters by cross-validated lasso regression. We adopt the second order Epanechnikov kernel as the weighting function $\omega$ of the continuous treatment, where the bandwidth corresponds to $2.34n^{-0.25}$. The latter choice corresponds to a Silverman (1986)-type rule of thumb for Epanechnikov kernels, as for instance also applied in Huber et al. (2020). In addition to estimation based on the rule of thumb, we consider an undersmoothed version, in which the kernel bandwidth is divided by 2. Furthermore, for estimating the generalized propensity scores $\rho_{d,t}(X)$ by lasso, we consider both linear and log-linear models that assume normally distributed unobservables $V$ (which is incorrect as $V$ is uniformly distributed). The ATET estimator also imposes a trimming rule that safeguards against disproportionately influential observations in the inverse propensity weighting (IPW) by dropping observations receiving a weight that is larger than 10% in the IPW-based computation of any conditional mean outcome given the treatment state and time period. Such a trimming rule for ensuring common support in the sample is considered in Huber et al. (2013). For obtaining the standard error of the ATET estimate, we consider the estimated asymptotic variance approximation provided in equation (20) of Section 4.
@@
The upper panel of Table 1 provides the simulation results for repeated cross-sections, when estimating the ATET $\Delta_{3,2,1}=E[Y_{1}(3)-Y_{0}(2)|D_{1}=3]=3^{2}-2^{2}=5$, based on linear lasso regression without undersmoothing (‘lasso’), loglinear lasso regression without undersmoothing (‘lnorm’), linear lasso regression with undersmoothing (‘under’), and loglinear lasso regression with undersmoothing (‘ln under’). The columns provide the estimation method (‘method’), the bias of the estimator (‘bias’), its standard deviation (‘std’), its root mean squared error (‘rmse’), the average standard error (‘avse’), and the coverage rate of the 95% confidence intervals (‘cover’), respectively, for either sample size.
@@
For $n=2000$, we observe that the ATET estimators without undersmoothing are non-negligibly biased, while those with undersmoothing are less biased at the price of having a somewhat higher variance. Overall, the undersmoothed versions provide a superior performance in terms of the root mean squared error (RMSE), which measures the overall estimation error as a combination of both bias and variance. Furthermore, the average standard error is generally close to the actual standard deviation, suggesting that asymptotic variance estimation as outlined in Section 4 works exceedingly well in our simulations. For the undersmoothed ATET estimators, this implies a reasonable coverage rate that is approximately the nominal level of 95%. Regarding the the estimators without undersmoothing, the coverage is notably low due to larger biases. The results are virtually not affected by the assumption of normal or lognormal models for the estimation of the generalized propensity scores. When increasing the sample size to $n=8000$, the biases of the estimators without undersmoothing are reduced, although still considerably different from zero. In contrast, the biases of the estimators with undersmoothing are negligibly small, being substantially better in terms of RMSE compared to the estimation without undersmoothing. In terms of coverage, the estimators without undersmoothing attain roughly 83%, while the undersmoothed versions attain 97%, which is nearly the nominal level.
@@
method bias std rmse avse cover bias std rmse avse cover
lasso -0.376 0.249 0.451 0.251 0.682 -0.173 0.182 0.252 0.183 0.843
lnorm -0.379 0.251 0.454 0.251 0.668 -0.176 0.182 0.253 0.184 0.840
under -0.090 0.283 0.297 0.287 0.949 -0.027 0.209 0.211 0.211 0.966
ln under -0.095 0.283 0.298 0.287 0.952 -0.031 0.210 0.212 0.211 0.967
@@
method bias std rmse avse cover bias std rmse avse cover
lasso -0.541 0.123 0.555 0.127 0.009 -0.220 0.056 0.227 0.055 0.023
lnorm -0.545 0.122 0.558 0.127 0.009 -0.224 0.056 0.231 0.055 0.021
under -0.067 0.103 0.123 0.111 0.921 0.039 0.058 0.070 0.061 0.910
ln under -0.071 0.103 0.125 0.111 0.916 0.034 0.058 0.067 0.060 0.922
Notes: columns ‘method’, ‘bias’, ‘std’, ‘rmse’, ‘avse’, and ‘cover’ provide estimation method, the bias of the estimator, its standard deviation, its root mean squared error, the average of the standard error, and the coverage rate of the 95% confidence intervals, respectively. The ‘lasso’ provides the results for lasso-based estimation when assuming a linear treatment model with normally distributed errors. The‘lnorm’ assumes instead a loglinear treatment model with normal errors. The ‘under’ relies on a linear treatment model with undersmoothing, using half of the bandwidth suggested by the rule of thumb for smoothing the continuous treatment. The ‘ln under’ provides the results for a loglinear treatment model and undersmoothing. The true ATET is 5.
@@
The lower panel of Table 1 provides the simulation results for the panel case, when again estimating the ATET $\Delta_{3,2,1}=5$. The biases of the ATET estimators without undersmoothing are once more nonnegligible under either sample size, while undersmoothing drives the bias to approach zero fast as the sample size increases, which entails a substantially superior performance in terms of RMSE. The average standard errors based on asymptotic variance approximation as outlined in Appendix 21 closely match the actual standard deviations of the respective estimators. Finally, we observe a slight under-coverage by the confidence intervals of the undersmoothed estimators. However, this is not substantially below the nominal level of 95%. In contrast, coverage is poor for the ATET estimators without undersmoothing, due to their large biases. For this reason, the simulation results for the panel and repeated cross-section cases are qualitatively fairly similar.
@@ 7 Conclusion
In this paper, we introduced the kernel-based treatment effect estimators for continuously distributed and possibly time-varying treatment doses in difference-in-differences (DiD) models. It adopts double machine learning to control for covariates and past treatment histories in a data-adaptive manner. We proposed methods for both repeated cross-sections, allowing covariate distributions within treatment groups to change over time (in contrast to most of the DiD literature), and panel data, along with asymptotic formulae for variance estimation. We also presented a simulation study, which indicates a decent finite sample performance of undersmoothed versions of our estimators in setups with several thousand observations. Our estimators are available in the causalweight for the statistical software R.
@@ A.1 Neyman orthogonality of the DR expressions
where the second equality follows from the law of iterated expectations. Therefore, considering Neyman-orthogonal functions for $\Psi_{d^{\prime},t^{\prime}}$ yields a Neyman-orthogonal expression of the ATET, which is an additive function of $\Psi_{d^{\prime},t^{\prime}}$ for $d^{\prime}$in $\{d,0\}$ and $t^{\prime}$in $\{t,0\}$.
@@
The Gateaux derivative of $E[\phi_{d^{\prime},t^{\prime}}]$ in the direction of $[\eta^{0}-\eta]$ is given by:
@@
As the first factor is zero due to the fact that $[\mu^{0}_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]=0$, the expression in the fourth line of equation (A.1) is also zero. Therefore, the Gateaux derivative of $E[\phi_{d^{\prime},t^{\prime}}]$ in the direction of $[\eta^{0}-\eta]$ is zero. Moreover, we trivially have that the Gateaux derivative of $E[Y_{T}|D=d,T=t]$ in the direction of $[\eta^{0}-\eta]$ is zero. We note that the DR expression (2) is equivalent to expression (A.1) when using the Neyman-orthogonal score functions (A.1) for computing $\Psi_{d,0}$, $\Psi_{0,t}$ $\Psi_{0,0}$, noticing that $E[Y_{T}|D=d,T=t]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot I\{T=t\}\cdot Y_{T}}{\Pi_{d,t}}\right]$. For this reason, the DR expression (2) satisfies the Neyman orthogonality.
@@
The Gateaux derivative of $E[\phi_{d_{t}^{\prime}}]$ in the direction of $[\eta^{0}-\eta]$ is given by:
@@
-- $\lim_{h\rightarrow 0}E\left[E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\right.\right.$
×$\times\left.\left.\frac{p_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)-p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}\bigg{|}\mathbf{D_{t-1}},X\right]\right]$
@@
Thus, the Gateaux derivative of $E[\phi_{d^{\prime}_{t}}]$ in the direction of $[\eta^{0}-\eta]$ is zero. Furthermore, the Gateaux derivative of $E[Y_{t}-Y_{t-1}|D_{t}=d_{t}]$ in the direction of $[\eta^{0}-\eta]$ is zero. We note that the DR expression (3) is equivalent to expression (A.1) while using the Neyman-orthogonal score functions (A.1) for computing $\Psi_{d^{\prime}_{t}}$, noting that $E[Y_{t}-Y_{t-1}|D_{t}=d_{t}]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot(Y_{t}-Y_{t-1})}{P_{d_{t}}}\right]$. As a result, the DR expression (3) satisfies the Neyman orthogonality.
@@
A.2 Sketch of proofs of asymptotic theories
The discussion below sketches the main ideas of the proofs for our asymptotic results, which closely follow Zhang [2024]. We focus on the panel data case and note that the arguments extend to the repeated cross-sections case with only minor changes. We recall that under the regularity conditions of Section 5, we defined the bias as the difference between the true ATET and its kernel counterpart when the continuous treatment intensity is approximated by a kernel function with bandwidth $h$:

### v2

@@ 7 Simulation study
In our simulations, we set the number of covariates to $p=100$ and the $j$th element in the coefficient vector $\beta$ to $0.4/i^{2}$ for $j\in\{1,...,p\}$, which implies a quadratic decay of covariate importance in terms of confounding. We consider two sample sizes of $n=2000,4000$. DiD estimation is based on the sample analog of equation (3). In our simulation design, there are only two periods $t$ such that $T=t$ and T=t1T=t-1 in equation (3) correspond to $T=1$ and $T=0$, respectively. Furthermore, there is no history of previous treatments in $T=0$ such that $\mathbf{D}_{t-1}=\emptyset$ in our simulations. However, from an econometric perspective, there is no difference between controlling for the treatment history $\mathbf{D}_{t-1}$ or covariates $X$, which are both time-varying sets of confounders. For this reason, not considering $\mathbf{D}_{t-1}$ in our simulations comes without loss of generality as we could easily relabel some elements in $X$ to represent $\mathbf{D}_{t-1}$.
For estimation, we apply the didcontDML function from the causalweight package by Bodory and Huber (2018) using the statistical software R (R Core Team (2020)). We implement this DML-based DiD estimator for repeated cross-sections with two-fold cross-fitting and estimate the nuisance parameters via cross-validated lasso regression. We employ a second-order Epanechnikov kernel as the weighting function $\omega$ for smoothing over the continuous treatment, with a bandwidth given by $2.34\cdot n^{-0.25}$. Choosing 2.34 as the bandwidth constant corresponds to a Silverman (1986)-type rule of thumb for second-order Epanechnikov kernels, as also applied, for instance, in Huber et al. (2020), while the rate $n^{-0.25}$ ensures undersmoothing to avoid asymptotic bias induced by kernel smoothing. In addition to estimation based on this rule, we consider a specification with stronger undersmoothing by multiplying the resulting bandwidth by 0.5 (which amounts to halving the constant in the bandwidth formula), yielding $2.34\cdot n^{-0.25}\cdot 0.5$.
@@
Furthermore, when estimating the generalized propensity scores $\rho_{d,t}(X)$ via lasso, we consider both linear and log-linear specifications that assume normally distributed unobservables $V$ (an assumption that is violated here, as $V$ is uniformly distributed). The ATET estimator additionally imposes a trimming rule to safeguard against disproportionately influential observations in the inverse propensity weighting (IPW). Specifically, observations receiving a weight larger than 10% in the IPW-based computation of any conditional mean outcome given the treatment state and time period are dropped. Such a trimming rule to ensure common support are also considered in Huber et al. (2013). Finally, standard errors for the ATET estimates are obtained using the estimated asymptotic variance approximation given in equation (26) of Section 5.
@@
The upper panel of Table 1 provides the simulation results for repeated cross-sections, when estimating the ATET $\Delta_{3,2,1}=E[Y_{1}(3)-Y_{0}(2)|D_{1}=3]=3^{2}-2^{2}=5$, based on linear lasso regression with weaker undersmoothing (‘lasso’), loglinear lasso regression with weaker undersmoothing (‘lnorm’), linear lasso regression with stronger undersmoothing (‘under’), and loglinear lasso regression with stronger undersmoothing (‘ln under’). The columns provide the estimation method (‘method’), the bias of the estimator (‘bias’), its standard deviation (‘std’), its root mean squared error (‘rmse’), and the average standard error (‘avse’), respectively, for either sample size.
@@
For $n=2000$, we observe that the ATET estimators with weaker undersmoothing are non-negligibly biased, whereas those with stronger undersmoothing are less biased while having a slightly higher standard deviation. Overall, the strongly undersmoothed versions have a superior performance in terms of a lower root mean squared error (RMSE), which measures the overall estimation error as the square root of the sum of the squared bias and variance. Moreover, the ATET estimates are very similar when using normal or lognormal models for the estimation of the generalized propensity scores. Concerning statistical inference, the average standard error (‘avse’) is generally close to the actual standard deviation (‘std’), suggesting that asymptotic variance estimation as outlined in Section 5 performs well in our simulations. When increasing the sample size to $n=8000$, the biases and standard deviations of all estimators decrease, but again, it is the strongly undersmoothed versions of the estimators that perform substantially better in terms of the RMSE than those with weaker undersmoothing.
@@
method bias std. dev. rmse avse bias std rmse avse
lasso -0.406 0.227 0.465 0.222 -0.202 0.152 0.253 0.151
lnorm -0.402 0.226 0.461 0.221 -0.199 0.151 0.250 0.150
under -0.107 0.241 0.263 0.241 -0.059 0.167 0.177 0.174
ln under -0.104 0.239 0.261 0.240 -0.057 0.167 0.177 0.173
@@
method bias std. dev. rmse avse bias std rmse avse
lasso -0.663 0.120 0.674 0.125 -0.358 0.055 0.362 0.053
lnorm -0.661 0.120 0.671 0.125 -0.357 0.055 0.361 0.053
under -0.177 0.100 0.203 0.106 -0.094 0.056 0.110 0.056
ln under -0.176 0.099 0.202 0.106 -0.094 0.056 0.109 0.056
Notes: columns ‘method’, ‘bias’, ‘std. dev.’, ‘rmse’, and ‘avse’ provide estimation method, the bias of the estimator, its standard deviation, its root mean squared error, and the average of the standard error, respectively. The ‘lasso’ provides the results for lasso-based estimation when assuming a linear treatment model with normally distributed errors. The‘lnorm’ assumes instead a loglinear treatment model with normal errors. The ‘under’ relies on a linear treatment model with strong undersmoothing, using half of the constant for bandwidth computation based on the rule of thumb for smoothing the continuous treatment. The ‘ln under’ provides the results for a loglinear treatment model and strong undersmoothing. The true ATET is 5.
@@
The lower panel of Table 1 provides the simulation results for the panel case when estimating the ATET $\Delta_{3,2,1}=5$. The absolute magnitudes of the biases and the RMSEs of the ATET estimators with weaker undersmoothing are again considerably larger that those of the strongly undersmoothed versions for either sample size. Moreover, the average standard errors based on asymptotic variance approximation for the panel case as outlined in Section 5 closely match the actual standard deviations of the respective estimators. For this reason, the simulation results for the panel and repeated cross-section cases are qualitatively similar.
8 Empirical application
In this section, we present an empirical application of our method in the context of the COVID-19 pandemic in Brazil—one of the countries most severely affected by the crisis. We estimate the causal effect of second-dose COVID-19 vaccination rates on mortality across Brazilian municipalities. Our data are compiled from multiple sources providing municipality-level health and socio-economic information, including the laboratory DB Molecular – Diagnósticos do Brasil, the Brazilian Institute of Geography and Statistics (IBGE), DataSUS (Ministry of Health, Brazil), and the monitoring of COVID-19 cases and deaths by Cota (2020). After cleaning and harmonizing the data, we obtain a balanced panel of 213 municipalities observed daily from August 4, 2020 to March 23, 2022, covering 597 days and yielding 127,161 observations.
We define the continuous and time-varying treatment variable as the municipality-specific vaccination rate, measured as the ratio of cumulative dose-2 vaccinations administered in a municipality up to a given day to the municipality’s population size according to the 2020 census. Figure 1 displays vaccination rates over time (across 597 days) during the sample period, highlighting considerable heterogeneity in the pace of vaccination across municipalities. Given the limited number of municipalities in our sample, we adopt a stacked design that permits pooling across multiple periods to increase statistical power. Specifically, we designate the first day on which any municipality reaches a $70\%$ (or 0.7) dose-2 vaccination rate as the reference day $t_{0}$, as indicated by the dashed vertical line in Figure 1. We then stack the data at seven-day intervals over five periods, $(t_{0},t_{7},t_{14},t_{21},t_{28})$, which increases the effective evaluation sample five-fold relative to a single day, yielding $5\cdot 213=1065$ panel units. Furthermore, this design avoids common support issues at the treatment intensities under comparison, which are chosen to be 60% (or 0.6) and 40% (or 0.4), as indicated by the horizontal dashed lines in Figure 1.
Figure 1: Dose-2 Vaccination Rate Over Time by Municipality
Our outcome variable is defined as the number of new COVID-19-related deaths per 100,000100{,}000 population in a given municipality at specific post-treatment horizons—namely one, two, four, seven, ten, fifteen, thirty, and sixty days after treatment. This allows us to study the dynamic evolution of the treatment effect. We control for the following covariates in a data-adaptive manner, conditional on which the parallel trends assumption is required to hold: (1) 45 demographic and socioeconomic characteristics measured at the municipal level, such as fertility, education, income, and life expectancy; (2) a set of dummy variables indicating the stacked reference days; (3) treatment histories capturing dose-2 vaccination rates one, two, seven, and fourteen days prior to the treatment day; and (4) dose-1 vaccination rates on the treatment day as well as their corresponding histories (dose-1 vaccination rates one, two, seven, and fourteen days prior to the treatment day), to avoid confounding from dose-1 vaccination.
Table 2 reports descriptive statistics (mean, standard deviation, minimum, and maximum) for several key variables - including the outcome, the treatment, and a selected set of covariates reflecting the dose-1 vaccination rate, socioeconomic conditions, and infrastructure - computed across the 213 municipalities and all 597 days of the balanced panel. While the mortality outcome and vaccination rates can vary at the daily level, municipalities’ socioeconomic and infrastructure characteristics are measured only once in 2022 and are therefore time-invariant (although they may exert time-varying influences on outcomes, which is why they are considered as potential control variables). The table shows that municipalities represent diverse socioeconomic contexts, with substantial variation in per capita income and infrastructure (such as sewerage network coverage and access to piped water), which are relevant factors for healthcare capacity and pandemic management. See Appendix A.3 for a more comprehensive set of descriptive statistics covering a broader range of variables.
Table 2: Descriptive statistics
variable mean std. dev. minimum maximum
COVID-19 deaths / population 0.0020 0.0014 0.0000 0.0075
dose-1 vaccination rate 0.26 0.30 0.00 1.30
dose-2 vaccination rate 0.36 0.35 0.00 1.07
municipal HDI 0.74 0.05 0.57 0.86
Gini index 0.50 0.05 0.35 0.63
per capita income 807.73 282.31 257.64 2043.74
proportion of poor (%) 8.69 8.08 0.44 43.46
sewage network (%) 68.49 28.91 0.71 99.83
piped water (%) 95.63 5.56 60.18 99.99
As in the simulation study in the previous section, we use the didcontDML function from the causalweight package by Bodory and Huber (2018) to estimate the causal effect of the vaccination rate, employing two-fold cross-fitting to mitigate overfitting. To flexibly adjust for high-dimensional control variables, the nuisance parameters are estimated nonparametrically using random forests as the machine learner (see, e.g., Ho, 1995; Breiman, 2001), assuming normally distributed unobservables in the generalized propensity score model. The kernel bandwidth used for smoothing over treatment intensity is set to $2.34\cdot n^{-0.25}\cdot 0.7$, that is, the rule-of-thumb constant for bandwidth selection multiplied by $0.7$ to implement undersmoothing based on $n^{-0.25}$. This choice lies between the weaker and stronger undersmoothing approaches considered in the simulation study. Finally, we apply a trimming rule that drops observations receiving more than 10% weight in the IPW-based estimation of the mean potential outcome under non-treatment.
Table 3 reports the estimated ATETs (measured in COVID-19 deaths per 100,000100{,}000 population) based on a comparison of a higher dose-2 vaccination rate of 60% ($d_{\text{treat}}=0.6$) versus a lower rate of 40% ($d_{\text{control}}=0.4$) across post-treatment horizons. The corresponding standard errors (SE) and p-values, clustered at the municipality level, are also reported. The results indicate that higher dose-2 coverage has no statistically significant short-run effects from one to fifteen days after treatment. However, statistically significant negative effects emerge at longer horizons, after thirty days (at the 5% level) and sixty days (at the 10% level). Specifically, the DML estimates suggest a reduction of approximately 1.5 COVID-19 deaths per 100,000100{,}000 inhabitants in municipalities with a 60% vaccination rate relative to a rate of 40%, after one to two months. These findings are consistent with the expected time lag between infection and death documented in the COVID-19 literature; see, e.g., Huber and Langen (2020).
Table 3: Effects of dose-2 on COVID-19 deaths with $d_{treat}=0.6$, $d_{control}=0.4$
post-treatment days ATET SE $p$-value
one day 0.057 0.355 0.874
two days 2.657 2.469 0.282
four days 2.798 2.543 0.271
seven days 2.962 4.437 0.504
ten days -0.072 2.733 0.979
fifteen days -4.352 3.077 0.157
thirty days -1.418 0.632 0.025
sixty days -1.583 0.824 0.055
@@ 9 Conclusion
In this paper, we have developed a generalized difference-in-differences framework to estimate the causal effects of time-varying continuous treatments. Our approach focused on identifying the Average Treatment Effect on the Treated (ATET) by comparing distinct treatment intensities, relying on a conditional parallel trends assumption that robustly accounts for complex treatment histories and covariate evolution. By integrating double/debiased machine learning with kernel-based weighting, we provided a rigorous estimation strategy that remains consistent even with high-dimensional nuisance parameters. Finally, we illustrated our method by applying it to Brazilian municipality-level data to estimate the effect of second-dose COVID-19 vaccination rates on mortality. The results indicate that higher vaccination coverage does not significantly affect mortality in the very short term but leads to a statistically significant reduction in COVID-19 deaths after several weeks, consistent with the expected time lag between infection and fatal outcomes.
@@ A.1 Neyman orthogonality of the DR expressions
where the second equality follows from the law of iterated expectations. Therefore, considering Neyman-orthogonal functions for $\Psi_{d^{\prime},t^{\prime}}$ yields a Neyman-orthogonal expression of the ATET, which is an additive function of $\Psi_{d^{\prime},t^{\prime}}$ for $d^{\prime}$ $\in$ $\{d,0\}$ and $t^{\prime}$ $\in$ $\{t,0\}$.
@@
The Gateaux derivative of $E[\phi_{d^{\prime},t^{\prime}}]$ in the direction of $\eta^{0}-\eta]$ is given by:
@@
As the first factor is zero due to the fact that $\mu^{0}_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]=0$, the expression in the fourth line of equation (A.1) is also zero. Therefore, the Gateaux derivative of $E[\phi_{d^{\prime},t^{\prime}}]$ in the direction of $\eta^{0}-\eta]$ is zero. Moreover, we trivially have that the Gateaux derivative of $E[Y_{T}|D=d,T=t]$ in the direction of $\eta^{0}-\eta]$ is zero. We note that the DR expression (2) is equivalent to expression (A.1) when using the Neyman-orthogonal score functions (A.1) for computing $\Psi_{d,0}$, $\Psi_{0,t}$ $\Psi_{0,0}$, noticing that $E[Y_{T}|D=d,T=t]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot I\{T=t\}\cdot Y_{T}}{\Pi_{d,t}}\right]$. For this reason, the DR expression (2) satisfies the Neyman orthogonality.
@@
The Gateaux derivative of $E[\phi_{d_{t}^{\prime}}]$ in the direction of $\eta^{0}-\eta]$ is given by:
@@
[…]
$\times\left.\left.\frac{p_{d^{\prime}_{t}}(\mathbf{D}_{t-1},X)-p^{0}_{d^{\prime}_{t}}(\mathbf{D}_{t-1},X)}{p^{0}_{d^{\prime}_{t}}(\mathbf{D}_{t-1},X)}\bigg|\mathbf{D}_{t-1},X\right]\right]$
@@
Thus, the Gateaux derivative of $E[\phi_{d^{\prime}_{t}}]$ in the direction of $\eta^{0}-\eta]$ is zero. Furthermore, the Gateaux derivative of $E[Y_{t}-Y_{t-1}|D_{t}=d_{t}]$ in the direction of $\eta^{0}-\eta]$ is zero. We note that the DR expression (4) is equivalent to expression (A.1) while using the Neyman-orthogonal score functions (A.1) for computing $\Psi_{d^{\prime}_{t}}$, noting that $E[Y_{t}-Y_{t-1}|D_{t}=d_{t}]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot(Y_{t}-Y_{t-1})}{P_{d_{t}}}\right]$. As a result, the DR expression (4) satisfies the Neyman orthogonality.
@@
A.2 Proofs of asymptotic theories
In this section, we provide the main ideas of the proofs for our asymptotic results and omit the details, which closely follow Chernozhukov et al. [2018], Chang [2020], and Zhang [2025]. We focus on the repeated cross-sectional case and note that the arguments for panel case follows from Zhang [2025] with minor notation changes.
First, we outline the main idea for Lemma 1. Recall that when introducing the regularity conditions in Section 6, we have defined the bias as the difference between the true ATET and its smoothed kernel counterpart when the continuous treatment intensity is approximated by a kernel function with bandwidth $h$: